OpenStudy (anonymous):

I am having trouble with this quadratic equation: x^2-x-2=0 I got: (x-2)(x+1)=0 x^2+ x - 2x - 2 x^2 - x - 2 x= 2 or x=-1 I am unsure, because the -1 does not work..

OpenStudy (anonymous):

your answer is correct man

OpenStudy (saifoo.khan):

Yeah, i will work!

OpenStudy (blacksteel):

-1 does work, actually. (-1)^2 -(-1) - 2 = 1 + 1 - 2 = 0.

OpenStudy (anonymous):

Yes, Blacksteel is correct